OpenStudy (anonymous):

limit of (sqrt(x+1) -1)/ x as x approaches 0.

OpenStudy (anonymous):

I think it would not exist.

OpenStudy (luigi0210):

\[\lim_{x \rightarrow 0}\frac{ \sqrt{x+1} -1}{ x }\]?

OpenStudy (jhannybean):

Would that mean you have to divide by the highest degree in the denominator?

OpenStudy (luigi0210):

If this is it then then multiply by the conjugate

OpenStudy (jhannybean):

Oh nevermind, Luigi is right.

OpenStudy (anonymous):

Okay. I was gonna do that but i thought it was just for approaching infinity.

OpenStudy (luigi0210):

Yea! for the first time in my life I'm right :P

OpenStudy (jhannybean):

\[\lim_{x \rightarrow 0}\frac{ \sqrt{x+1} }{ x }*\frac{ \sqrt{x-1} }{ \sqrt{x-1} }\]

OpenStudy (anonymous):

would it be 1?

OpenStudy (luigi0210):

and if you do it right then you should get an answer of 1/2

OpenStudy (luigi0210):

Bean you forgot the -1 D:

OpenStudy (jhannybean):

No, wouldnt be 1 i don't think. Depends on how you solve it?

OpenStudy (anonymous):

yea! i know what i did wrong

OpenStudy (jhannybean):

Luigi there is no edit option D:

OpenStudy (luigi0210):

Haha that sucks. But anyways lucy do you understand?

sam (.sam.):

Are you familiar with L'Hospital's or whatever that is

OpenStudy (anonymous):

Yea i understand. but wouldn't the answer be DNE?

OpenStudy (anonymous):

OMG! Nevermind!

OpenStudy (luigi0210):

You don't really need that here Sam.. I think

OpenStudy (luigi0210):

and no bean forgot that -1 at the end *cough*

OpenStudy (jhannybean):

LH rule applies when the limit goes to infinity and becomes indeterminant, correct?

OpenStudy (anonymous):

it is 1/2, just wrote - instead of +

OpenStudy (jhannybean):

<_<

OpenStudy (luigi0210):

good job :P

OpenStudy (anonymous):

I have one more i have trouble with. could you help me?

OpenStudy (jhannybean):

did you finish solving it?

OpenStudy (luigi0210):

and I gtg, so @Jhannybean has got you from here. Good luck! :P

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x ^{2} } - \frac{ 5x }{ x + 2 }\] as x approaches infinity.

OpenStudy (anonymous):

I am not sure if i would have to have a common denominator or if it would be undefined.

OpenStudy (jhannybean):

I am going to make my best attempt at solving this problem :P

OpenStudy (jhannybean):

\[\lim_{x \rightarrow \infty}((\frac{ 1 }{ 3x^2 })-\frac{ 5x }{ x+2 })\] You can split the limit and solve each one separately\[\lim_{x \rightarrow \infty}\frac{ 1 }{ 3x^2 }- \lim_{x \rightarrow \infty}\frac{ 5x }{ x+2 }\]

OpenStudy (anonymous):

the first would be negligible and the second would go to 5? so -5?

OpenStudy (jhannybean):

\[\frac{ 1 }{ 3 }\lim_{x \rightarrow \infty}x^2 - \lim_{x \rightarrow \infty}\frac{ \frac{ 5 }{ x } }{ \frac{ (x+2) }{x } }\]

OpenStudy (jhannybean):

oops i forgot the x of 5x! yes it would be 5

OpenStudy (jhannybean):

(5x/x)/[(x+2)/x]

OpenStudy (anonymous):

\[\frac{ \sqrt{x+1} -1}{ x }\times \frac{\sqrt{x+1}+1}{\sqrt{x+1}+1}\] \[\frac{x}{x(\sqrt{x+1}+1)}\] \[=\frac{1}{\sqrt{x+1}+1}\]

OpenStudy (anonymous):

thanks! @Jhannybean

OpenStudy (jhannybean):

No problem! and Jim helped answer your previous question :P

OpenStudy (jhannybean):

wow i mean @satellite73